I am having trouble solving the system of equations by graphing. I think I have it and then I am off by one number.
Here is an example
x-2y=8
x+2y=-4
I need to find the solution set and graph.

Thank you:)

Answer by mananth(16946)

1 x -2 y = 8
1 x + 2 y = -4

equation I
when x= 0 y= -4 ,( 0 , -4 )
when y = 0 x= 8 ,( 8 , 0 )
Equation 2
when x= 0 y= -2 ( 0 , -2 )
when y = 0 x= -4 ( 0 , -4 )



You will observe that the point of intersection is 2,-3 The solution of the system is ( 2,-3 )